# 7491, Trondheim, Norway

### v1

@@ 6.2 Problems governed by quasi-convex functions
If all $f_{i}$ are convex, we have the same conclusions in Theorem 6.1 for the solution map $\Theta_{1}$ defined by (5.1). However, until now, we have not known any explicit convex vector optimization problems whose (weak) Pareto solution set is disconnected. This is related to Question 9.8 in yen2011monotone and Question 5.1 in hieu2021disconnectedness .
In the case that all objective functions are strictly quasi-convex, $\Theta_{2}$ is single-valued on its domain. Therefore, we have the following theorem:
@@
Suppose that all $f_{i}$ are strictly quasi-convex and strictly bounded from below by 00 on $K$. Then, one has $\operatorname{Sol}(K,f)=\operatorname{Sol^{w}}(K,f)$. Moreover, if $\operatorname{Sol}(K,f)$ is disconnected then all its components are unbounded.
@@
For $\xi\in\Delta$, $f^{\xi}$ is also strictly quasi-convex on $K$ (see (warburton1983quasiconcave, , Lemma 4.2)); Hence $\Theta_{2}$ is single-valued on its domain. The conclusions are implied straightforwardly from Corollary 1. ∎
@@
The results related to set-valued mappings having connected values in this paper can be useful for investigating (un)bounedness and (un)connectedness of solution sets in set optimization, vector equilibrium, or vector variational inequality under some generalized convexity/monotone assumptions. We plan to exploit special structures such as polyhedral or semi-algebraic to discover new properties concerning contractibility or connectedness of solution sets of these problems as in huong2017connectedness ; huong2016polynomial ; hieu2019polynomialVVIS ; hieu2020application .

### v2

@@ 6.3 Problems governed by convex functions
If $K$ is convex and all $f_{i}$ are convex over $K$, we have the same conclusions in Theorem 6.1 for the solution map $\Theta_{1}$ defined by (5.1). However, until now, we still do not know any explicit convex vector optimization problems whose (weak) Pareto solution set is disconnected. This is related to Question 9.8 in yen2011monotone and Question 5.1 in hieu2021disconnectedness .
When $\operatorname{VOP}(K,f)$ is given by convex polynomial data, i.e. $f_{1},\dots,f_{m}$ are polynomials and $K$ is given by convex polynomials as follows:
$K=\{x\in\operatorname{\mathbb{R}}^{n}:g_{1}(x)\leq 0,\dots,g_{s}(x)\leq 0\},$
under a mild condition, in the following theorem, we claim that $\operatorname{Sol^{w}}(K,f)$ is connected.
@@
Let $f_{1},\dots,f_{m}$ and $g_{1},\dots,g_{s}$ be polynomial and convex. Assume that $\Theta_{1}$ has bounded values. Then, the weak Pareto solution set is connected.
@@
Recall that $f_{\xi}$ is also convex over $K$ and $\operatorname{Sol^{w}}(K,f)=\Phi_{1}(\Delta)$ (see Lemma 4). Assume that $\operatorname{Sol^{w}}(K,f)$ is nonempty, this implies that $\operatorname{dom}\Phi_{1}$ is so. We prove that $\operatorname{dom}\Phi_{1}$ is convex. Indeed, let $\xi^{\prime},\xi^{\prime\prime}$ be two points in $\operatorname{dom}\Phi_{1}$, i.e., $f_{\xi^{\prime}}$ and $f_{\xi^{\prime\prime}}$ are bounded from below over $K$. This implies that $f_{\xi}$ is also bounded from below over $K$ for all $\xi$ in the line segment connecting $\xi^{\prime}$ and $\xi^{\prime\prime}$, denoted by $[\xi^{\prime},\xi^{\prime\prime}]$. According to the Frank–Wolfe type theorem in convex polynomial optimization belousov2002frank , $\operatorname{Sol}(K,f_{\xi})$ is nonempty for all $\xi\in[\xi^{\prime},\xi^{\prime\prime}]$. Hence, $\operatorname{dom}\Phi_{1}$ is convex.
Therefore, connectedness of $\operatorname{Sol^{w}}(K,f)$ is straightforward from Lemma 5 and Theorem 2.1. ∎
To end this section, we illustrate Theorems 6.3 and 6.4 by considering the following example:
Example 8
Consider the convex polynomial bicriteria optimization problem $\operatorname{VOP}(K,f)$ in $\operatorname{\mathbb{R}}^{2}$, with the convex feasible set $K=\operatorname{\mathbb{R}}_{+}^{2}$ and two convex objective functions given by
$f_{1}(x)=x_{1}^{2}-x_{2},\quad f_{2}(x)=-x_{1}+x_{2}^{2}.$
Clearly, these functions are not bounded from below on $K$. Thus, the two end points $(0,1),(1,0)$ in $\Delta$ do not belong to the domain of the solution mapping $\Theta_{1}$ of $\operatorname{VOP}(K,f)$. Let $\xi\in\operatorname{ri}\Delta$, equivalently $\xi_{1}\in(0,1)$. One has
$f_{\xi}=\xi_{1}f_{1}+(1-\xi_{1})f_{2}=\xi_{1}x_{1}^{2}+(1-\xi_{1})x_{2}^{2}-(1-\xi_{1})x_{1}-\xi_{1}x_{2},$
which is quadratic and convex over $K$. An easy computation shows that $\operatorname{Sol}(K,f_{\xi})$ has the unique solution
$\operatorname{Sol}(K,f_{\xi})=\Big{\{}\Big{(}\frac{1-\xi_{1}}{2\xi_{1}},\frac{\xi_{1}}{2(1-\xi_{1})}\Big{)}\Big{\}},\,\xi_{1}\in(0,1).$
One sees that the weak Pareto solution set is connected, moreover, the set is unbounded and $\Theta_{1}$ has non-closed domain.
@@
The results related to set-valued mappings having connected values in this paper can be useful for investigating (un)bounedness and (dis)connectedness of solution sets in set optimization, vector equilibrium, or vector variational inequality under some generalized convexity/monotone assumptions. We plan to exploit special structures in geometry such as polyhedral or semi-algebraic to discover new properties concerning contractibility or connectedness of solution sets of these problems as in huong2017connectedness ; huong2016polynomial ; hieu2019polynomialVVIS ; hieu2020application .
Acknowledgements.
The authors are indebted to the anonymous referee for the insightful comments and valuable suggestions. The ERCIM Alain Bensoussan Fellowship Programme supported the research of Vu Trung Hieu.